Write each equation in standard form using integers.
y=2/3x+7y
(the 2/3x is a fraction)

Answer by jim_thompson5910(35256)

y=2/3x+7y


3y=2x+21y


-2x+3y=21y


-2x+3y-21y=0


-2x-18y=0


2x+18y=0


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim